1.             What mass of calcium sulfate can be produced from the reaction of 1250 g calcium phosphate with 750 g H2SO4?

 

Ca3(PO4) 2 (s) +  H2SO4(aq) à   CaSO4(s) + H3PO4(aq)


Give the given, unknown and full solution.

1
Expert's answer
2022-04-05T10:35:05-0400

Moles= mass/ rfm

Moles of calcium phosphate = 1250/310= 4moles

Moles of H2SO4= 750/98= 7.7moles

Mole ratio= 1:1

Mass of calcium sulfate= 4× 136= 544g